OpenStudy (anonymous):

a recipe calls for 5/8cup butter . angie wants to triple the recipie but she has only 1 5/8 cups of butter . how much butter does she need to borrow

jimthompson5910 (jim_thompson5910):

5/8 times 3 = ???

jimthompson5910 (jim_thompson5910):

what do you get?

OpenStudy (anonymous):

15/8

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

so she needs 15/8 cups 15/8 = (8+7)/8 15/8 = 8/8+7/8 15/8 = 1+7/8 15/8 = 1 7/8 So she needs 1 7/8 cups

jimthompson5910 (jim_thompson5910):

Now subtract